OpenStudy (loser66):

A population of ants, initially containing 20 ants, is growing according to the population growth equation \(P'(t) =0.5P(t)(1-\dfrac{P(t)}{N})\) with time t measured in days. 1) If the carrying capacity of the area in which they live is 1000, how many ants will there be one week later? 2) If bug spray is used, and it decreases the intrinsic growth rate r from 0.5 to 0.1, how many ants will you have one week later? Assume the carrying capacity remains unchanged at 1000 ants. Please, help

OpenStudy (loser66):

@ikram002p HI, friend. For some reasons, I got the wrong answer but don't know how to fix. Help me, please.

OpenStudy (loser66):

and the most frustrating thing is it is just the first order of ODE, no tricky. It takes me down !!!

OpenStudy (sparklestaraa):

Screenshot the question for me

OpenStudy (loser66):

Why? it is what I type.

OpenStudy (loser66):

ok, I will.

OpenStudy (loser66):

oh, you mean my work?

OpenStudy (sparklestaraa):

No the question because this "\(P'(t) =0.5P(t)(1-\dfrac{P(t)}{N})\) " is not clear

OpenStudy (sparklestaraa):

Oh never mind sorry

OpenStudy (loser66):

Ok, I show you what I got.

OpenStudy (sparklestaraa):

Ok

OpenStudy (loser66):

\(\dfrac{dP}{dt}=0.5P\dfrac{N-P}{N}\) \(\dfrac{dP}{P(N-P)}=\dfrac{0.5}{N}dt\) integral both sides

OpenStudy (loser66):

Decompose \(\dfrac{1}{P(N-P)}=\dfrac{1}{P}+\dfrac{1}{N-P}\) Hence \(ln|\dfrac{P}{N-P}|= \dfrac{0.5t}{N}+C\)

OpenStudy (loser66):

That gives us \(\dfrac{P}{N-P}=Ae^{0.5t/N}\) Is there any mistake?

OpenStudy (loser66):

Now, using initial value P(0) =20 to find A, with N =1000 \(\dfrac{20}{1000-20}= A = 0.0204\)

OpenStudy (loser66):

Hence general solution is \(P = 0.00204e^{0.00005t}(1000-P)\) Hence \(P=\dfrac{0.00204e^{0.00005t}}{1+0.00204e^{0.00005t}}\) But if there is no mistake , then P(7) = 0.00203 which is non sense.

OpenStudy (loser66):

oh, I forgot *1000 but even though , P(7) = 2.03 is nonsense till. :(

OpenStudy (freckles):

I think your partial fraction decomposition is incorrect

OpenStudy (freckles):

\[\frac{1}{P(N-P)}=\frac{a}{P}+\frac{b}{N-P} \\ 1=a(N-P)+bP \\ 1=aN+P(-a+b) \\ \implies 1=a N \text{ and } -a+b=0\]

OpenStudy (freckles):

\[\frac{1}{P(N-P)}=\frac{\frac{1}{N}}{P}+\frac{\frac{1}{N}}{N-P}\]

OpenStudy (freckles):

N is carrying capacity?

OpenStudy (freckles):

I'm getting a answer for P(7) in the 4 hundreds

OpenStudy (freckles):

\[P=\frac{1000 e^{.5 t}}{49+e^{.5t }}\] is what I have for P if N is 1000

OpenStudy (freckles):

oh and it looks pretty like this because I left a in the prettier form

OpenStudy (freckles):

fraction form that is (not approximated)

OpenStudy (loser66):

Ok, let me redo. \(\dfrac{dP}{dt}= 0.5P(1-\dfrac{P}{N})\) \(\dfrac{dP}{P(1-\dfrac{P}{N})}=0.5dt\) Now decompose the LHS \(\dfrac{a}{P}+\dfrac{b}{1-P/N}=1\) that gives us \((1-P/N) a+bP=1\) 1) if P =0, we have a=1 2) if P =N , we have bN =1or b =1/N replace all \(\int \dfrac{1}{P}dP +\int\dfrac{1/N}{1-P/N} dP\) simplify more for the second term, we have \[\int \dfrac{dP}{P}-\int\dfrac{dP}{P-N}= ln|\dfrac{P}{P-N}|= 0.5t +C\]

OpenStudy (loser66):

That gives \(\dfrac{P}{P-N}=Ae^{0.5t}\) If replace P(0) =20, we get \(\dfrac{20}{20-1000}=A = -0.0204\)

OpenStudy (loser66):

And then solve for P. I got \(P=Ae^{0.5t}(P-N)\rightarrow P=\dfrac{-ANe^{0.5t}}{1-Ae^{0.5t}}\)

OpenStudy (freckles):

\[P'=\frac{1}{2}P(\frac{N-P}{N}) \\ \frac{P'}{P(N-P)}=\frac{1}{2} \frac{1}{N} \\ \int\limits \frac{dP}{P(N-P)} dP = \int\limits \frac{1}{2} \frac{1}{N} dt \\ \text{ \partial fraction time!} \\ \frac{1}{P(N-P)}= \frac{a}{P}+\frac{b}{N-P} \\ 1=a(N-P)+b(P) \\ 1=aN+P(-a+b) \\ -a+b=0 \implies a=b \\ 1=a N \implies a=b=\frac{1}{N} \\ \\ \text{ so we have after the integration part } \\ \frac{1}{N}(\ln(P)-\ln(N-P))=\frac{1}{2N} t +C \\ \ln(\frac{P}{N-P})=\frac{1}{2}t +k \\ \frac{P}{N-P}=a e^{\frac{1}{2} t} \\ P=(N-P)a e^{\frac{1}{2}t} \\ P=N ae^{\frac{1}{2}t}-Pae^{\frac{1}{2}t} \\ P(1+ae ^{\frac{1}{2}t})=Na e^{\frac{1}{2}t} \\ P=\frac{N a e^{\frac{1}{2}t}}{1+ae^{\frac{1}{2}t} } \text{ where I have } A=\frac{1}{49} \\ P=\frac{1000 \cdot \frac{1}{49} e^{\frac{1}{2}t}}{1+\frac{1}{49} e^{\frac{1}{2}t}}\] so yeah your P looks equivalent to mine

OpenStudy (freckles):

because when you plug in your values you get that same P

OpenStudy (freckles):

and I just multiplied numerator and denominator by 49 to make it look prettier

OpenStudy (freckles):

\[P=\frac{1000 e^{\frac{1}{2}t}}{49+e^{\frac{1}{2}t}}\]

OpenStudy (freckles):

so you should give P(7) equals some number in the 4 hundreds now

OpenStudy (loser66):

If a= b= 1/N, then plug in we have \[\int \dfrac{1/N}{P}+\dfrac{1/N}{N-P}=1/N \int \dfrac{1}{P}-\dfrac{1}{P-N}=1/Nln|\dfrac{P}{P-N}|\] You see, my logic gives me the denominator is P-N while yours is N-P. It makes our solutions different.

OpenStudy (loser66):

eventhoug |P-N|=|N-P| but I want to make it clear when we take off the absolute value.

OpenStudy (freckles):

?

OpenStudy (freckles):

\[\text{ you got } A=\frac{-1}{49} \\ \text{ and your } P \text{ was } \\ P=\frac{-ANe^{\frac{1}{2}t}}{1-Ae^{\frac{1}{2}t}} \\ \\ \text{ also we have } N=1000 \text{ plug \in we have } \\ P=\frac{-(\frac{-1}{49}) (1000)e^{\frac{1}{2}t}}{1-\frac{-1}{49} e^{\frac{1}{2}t}} =\frac{\frac{1000}{49}e^{\frac{1}{2}t}}{1+\frac{1}{49}e^{\frac{1}{2}t}} =\frac{1000e^{\frac{1}{2}t}}{49+e^{\frac{1}{2}t}}\] this is exactly the same I have abov

OpenStudy (freckles):

our solutions are exactly the same

OpenStudy (freckles):

So I'm not really sure what you are talking about